shift summary for owl 6/30/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Wed, 30 Jun 1999 07:04:14 -0500

Shift Summary for owl shift, 30 Jun 99

On Shift: Leo Bellantoni, Ashkan Alavi

Hours of beam delivered at all intensities: 7.0
Hours of beam delivered at nominal intensity: 7.0
Hours of E832 data (special runs not included): 5.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Various DAQ problems.

========================================================================
Runs ended this shift:
Run 13766 - 1.0M events - MUON
Run 13767 - 1.2M events - BE~AM_832 sem = 5.0E12
Run 13772 - In progress - BEAM_832 sem = 5.6E12

Tapes brought to FCC this shift: QKC061-66, QKK061-66, QKS061-66,
QKB021-22

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:

Finished the muon runs from the last shift. Ran BEAM_832 at sem=5E12 for
the rest of the shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13766 00:04 Muon run at nominal DC voltage starts.
00:25 Run ends. Ask for hadrons. Analysis magnet
polarity flipped to negative. At OSPLIT request,
pop tape QKMU08 into drive...

13767 00:47 Run begins.
01:15 Run aborts! Many error messages: "DFN provider is
not currently active". Problem was failure of
tape drive 2 of KTeV3; Ed fixed the destination
logic remotely; we drt_fresh_start. The new
destinations use TPS2D3 of KTEV3, which is still
probably on the list of tape drives allocated for
the online split => POTENTIAL FUTURE PROBLEM~
13768 02:16 DAQ is running again, but front end isn't. At
various times, KTeVana errors from many different
parts of the readout, stream 5 permit gone, or
both streams 5 & 6 not getting permit back. Tried
cycling repeater downstairs, re-initing KQs, BANs,
pipeline...
13772 1 03:38 Running finally. Cure was complete
drt_fresh_start; reason why this helped far from
obvious. Perhaps the trigger was not correctly
initialized?
13772 04:10 Gas checklist: All Ar-Eth cylinders outside the
shed are empty. Inside the shed, there is plenty.
Email fired off to appropriate parties.
13772 05:00 Coffee supplies dangerously low. And not much
duck tape left, either.
13722 05:20 COSMic counter histos look poor. Email sent to
appropriate parties.
13722 156 06:55 Laser RMS too wide, but it goes away in next
spill. It was channel 1451.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Tape drive 3 on Ktev1 was configured incorrectly. Needs to be
reconfigured in hardware tomorrow dayshift when we have a down time.

Tape drive tps2d3 of Ktev3 is being used for QKB output, but is
probably still listed as an available drive by split software.

Coffee situation critical.